11 11 TEX L^{A}TEX

The Linux HOWTO Index

by Guylhem Aznar, guylhem@oeil.qc.ca v3, February 2000

This document contains an index to the Linux HOWTOs as well as other information about the HOWTO project.

Contents

1 What Are Linux HOWTOs? 1

2 Where Can I Get Linux HOWTOs? 1 2.1 HOWTO Translations . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

3 Index 2 3.1 HOWTOs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3.2 Special HOWTOs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 3.3 Unmaintained HOWTOs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

4 Writing and Submitting a HOWTO 11

5 Copyright 12

1 What Are Linux HOWTOs?

Linux HOWTOs are documents which describe in detail a certain aspect of configuring or using Linux. For example, there is the Installation HOWTO, which gives instructions on installing Linux, and the Mail HOWTO, which describes how to set up and configure mail under Linux. Other examples include the NET-3 HOWTO and the Printing HOWTO. HOWTOs are comprehensive docs - much like an FAQ but generally not in question-and-answer format. However, many HOWTOs contain an FAQ section at the end. There are several HOWTO formats available: plain text, PostScript, DVI, and HTML. In addition to the HOWTOs, there are a multitude of mini-HOWTOs on short, specific subjects.

2 Where Can I Get Linux HOWTOs?

HOWTOs can be retrieved via anonymous FTP from the following sites:

   • ftp://metalab.unc.edu/pub/Linux/docs/HOWTO
   • ftp://tsx-11.mit.edu/pub/linux/docs/HOWTO

as well as many mirror sites. You can also browse HOWTOs in HTML format <http://metalab.unc.edu/LDP/HOWTO/> or grab All-html-

HOWTOs.tar.gz <http://metalab.unc.edu/LDP/HOWTO/All-html-HOWTOs.tar.gz> for viewing on your

own PC. The files will be extracted into HOWTO/html/.

Many mirror sites <http://metalab.unc.edu/LDP/hmirrors.html> mirror the HTML files.

metalab.unc.edu is heavily used, so please use a mirror site if possible. HOWTOs are also posted towards the beginning of the month to the Usenet newsgroup

comp.os.linux.answers. There is a tool called NewstoHOWTO <ftp://metalab.unc.edu/pub/Linux/

system/news/misc/> that will assemble the postings.

2.1 HOWTO Translations

HOWTO translations are available on metalab.unc.edu <ftp://metalab.unc.edu/pub/Linux/docs/HOWTO/ translations> and mirrors around the world. So far there are:

• Chinese (zh)

• Croatian (hr)

• French (fr)

• German (de)

• Hellenic (el)

• Indonesian (id)

• Italian (it)

• Japanese (ja)

• Korean (ko)

• Polish (pl)

• Slovenian (sl)

• Spanish (es)

• Swedish (sv)

• Turkish (tr) If you know of any other translation projects, please let me know and I will add them to this list. If you are interested in getting your translations archived on metalab.unc.edu, please read the directory structure specification at http://metalab.unc.edu/pub/Linux/docs/HOWTO/translations/Directory-Structure and get in touch with me.

3 Index

3.1 HOWTOs

The following Linux HOWTOs are currently available:

• 3Dfx-HOWTO, updated: February 1998,

• AI-Alife-HOWTO, updated: Jan 2000,

• AX25-HOWTO, updated: October 1997,

• Access-HOWTO, updated: March 1997,

• Alpha-HOWTO, updated: June 1997,

• Assembly-HOWTO, updated: February 2000,

• Bash-Prompt-HOWTO, updated: December 1999,

• Belgian-HOWTO, updated: September 1999,

• Benchmarking-HOWTO, updated: August 1997,

• Beowulf-HOWTO, updated: November 1998,

• BootPrompt-HOWTO, updated: May 1999,

• Bootdisk-HOWTO, updated: January 2000,

• Busmouse-HOWTO, updated: June 1998, • C-C++Beautifier-HOWTO, updated: February 2000 ,

• CD-Writing-HOWTO, updated: September 1999,

• CDROM-HOWTO, updated: March 1999,

• CVS-HOWTO, updated: December 1999,

• CVS-RCS-HOWTO, updated: November 1999,

• Chinese-HOWTO, updated: June 1998,

• Commercial-HOWTO, updated: March 1999,

• Config-HOWTO, updated: September 1999,

• Consultants-HOWTO, updated: September 1999,

• Cyrillic-HOWTO, updated: January 1998,

• DNS-HOWTO, updated: February 1999,

• DOS-Win-to-Linux-HOWTO, updated: February 1999,

• DOSEMU-HOWTO, updated: March 1997,

• Danish-HOWTO, updated: February 1999,

• Diskless-HOWTO, updated: Jan 2000,

• Diskless-root-NFS-HOWTO, updated: March 1999,

• Distribution-HOWTO, updated: February 2000,

• Ecology-HOWTO, updated: January 2000,

• Emacs-Beginner-HOWTO, updated: October 1999,

• Emacspeak-HOWTO, updated: April 1999,

• Enterprise-Java-for-Linux-HOWTO, updated: January 1999,

• Esperanto-HOWTO, updated: August 1999,

• Ethernet-HOWTO, updated: May 1999,

• Finnish-HOWTO, updated: February 1996,

• Firewall-HOWTO, updated: September 1999,

• Font-HOWTO, updated:

• Framebuffer-HOWTO, updated: July 1999,

• Francophones-HOWTO, updated: Décembre 1999,

• From-PowerUp-To-Bash-Prompt-HOWTO, updated: February 2000,

• Ftape-HOWTO, updated: August 1998, • GCC-HOWTO, updated: February 1996,

• German-HOWTO, updated: March 1997,

• Glibc2-HOWTO, updated: June 1998,

• HOWTO-HOWTO, updated: February 1999,

• HOWTO-INDEX, updated: February 2000,

• Hardware-HOWTO, updated: September 1999,

• Hebrew-HOWTO, updated: September 1995,

• Hellenic-HOWTO, updated: Augoustos 1997,

• INFO-SHEET, updated: September 1998,

• IP-Masquerade-HOWTO, updated: January 2000,

• IPCHAINS-HOWTO, updated: March 1999,

• IPX-HOWTO, updated: May 1998,

• IR-HOWTO, updated: September 1999,

• ISP-Hookup-HOWTO, updated: March 1998,

• IngresII-HOWTO, updated: December 1999,

• Installation-HOWTO, updated: November 1999,

• Intranet-Server-HOWTO, updated: August 1997,

• Italian-HOWTO, updated: November 1998,

• Java-CGI-HOWTO, updated: December 1998,

• Jaz-Drive-HOWTO, updated: January 2000,

• Kernel-HOWTO, updated: June 1999,

• Keyboard-and-Console-HOWTO, updated: February 1998,

• KickStart-HOWTO, updated: January 1999,

• Kiosk-HOWTO, updated: October 1999,

• LDAP-HOWTO, updated: June 1999,

• LFS-HOWTO, updated: February 2000,

• Laptop-HOWTO, updated: December 1999,

• Large-Disk-HOWTO, updated: November 1999,

• Linux-From-Scratch-HOWTO, updated: January 2000,

• LinuxDoc+Emacs+Ispell-HOWTO, updated: February 1998, • Loopback-Encrypted-Filesystem-HOWTO, updated: November 1999,

• META-FAQ, updated: October 1997,

• MGR-HOWTO, updated: May 1996,

• MILO-HOWTO, updated: December 1996,

• MIPS-HOWTO, updated: March 1999,

• MP3-HOWTO, updated: September 1999,

• Mail-Administrator-HOWTO, updated: January 2000,

• Mail-User-HOWTO, updated: May 1999,

• Majordomo-MajorCool-HOWTO, updated: November 1999,

• Modem-HOWTO, updated: January 2000,

• Multi-Disk-HOWTO, updated: November 1999,

• Multicast-HOWTO, updated: March 1998,

• Mutt-GnuPG-PGP-HOWTO, updated: February 2000,

• NC-HOWTO, updated: December 1999,

• NCD-HOWTO, updated: Sept 1999,

• NET3-4-HOWTO, updated: August 1999,

• NFS-HOWTO, updated: October 1999,

• NIS-HOWTO, updated: March 1999,

• Net-HOWTO, updated: December 1999,

• Networking-Overview-HOWTO, updated: September 1999,

• Online-Troubleshooting-HOWTO, updated: September 1999,

• Optical-Disk-HOWTO, updated: December 1998,

• Oracle-HOWTO, updated: August 1998,

• PCI-HOWTO, updated: March 1997,

• PCMCIA-HOWTO, updated: December 1999,

• PLIP-Install-HOWTO, updated: November 1999,

• PPP-HOWTO, updated: March 1997,

• PalmOS-HOWTO, updated: September 1998,

• Parallel-Processing-HOWTO, updated: January 1998,

• Plug-and-Play-HOWTO, updated: January 2000, • Polish-HOWTO, updated: February 1999,

• Portuguese-HOWTO, updated: de 1999,

• PostgreSQL-HOWTO, updated: December 1999,

• Printing-HOWTO, updated: Feb 2000,

• Printing-Usage-HOWTO, updated: February 1998,

• Process-Monitor-HOWTO, updated: December 1999,

• Psion-HOWTO, updated: Nov 1999,

• Quake-HOWTO, updated: August 1998,

• Reading-List-HOWTO, updated: November 1999,

• Root-RAID-HOWTO, updated: March 1998,

• SCSI-Programming-HOWTO, updated: May 1996,

• SMB-HOWTO, updated: June 1999,

• SMP-HOWTO, updated: January 2000,

• SRM-HOWTO, updated: December 1999,

• Secure-Programs-HOWTO, updated: February 2000,

• Securing-Domain-HOWTO, updated: and 2049,

• Security-HOWTO, updated: April 1999,

• Serbian-HOWTO, updated: November 1998,

• Serial-HOWTO, updated: January 2000,

• Serial-Programming-HOWTO, updated: January 1998,

• Shadow-Password-HOWTO, updated: April 1996,

• Slovenian-HOWTO, updated: February 1999,

• Software-Building-HOWTO, updated: July 1999,

• Software-RAID-HOWTO, updated: November 1998,

• Software-Release-Practice-HOWTO, updated: January 2000,

• Sound-HOWTO, updated: March 1999,

• Sound-Playing-HOWTO, updated: August 1998,

• Spanish-HOWTO, updated: August 1996,

• TclTk-HOWTO, updated: November 1998,

• TeTeX-HOWTO, updated: November 1998, • Text-Terminal-HOWTO, updated: December 1999,

• Thai-HOWTO, updated: August 1998,

• Thinclient-HOWTO, updated: November 1999,

• Tips-HOWTO, updated: June 1998,

• Turkish-HOWTO, updated: September 1999,

• UMSDOS-HOWTO, updated: November 1995,

• UPS-HOWTO, updated: November 1997,

• UUCP-HOWTO, updated: February 1998,

• Unicode-HOWTO, updated: October 1999,

• Unix-Internet-Fundamentals-HOWTO, updated: September 1999,

• Unix-and-Internet-Fundamentals-HOWTO, updated: February 2000,

• User-Group-HOWTO, updated: April 1998,

• VAR-HOWTO, updated: August 1999,

• VME-HOWTO, updated: July 1998,

• VMS-to-Linux-HOWTO, updated: September 1999,

• VPN-HOWTO, updated: Dec 1999,

• VPN-Masquerade-HOWTO, updated: September 1999,

• Vim-HOWTO, updated: December 1999,

• Virtual-Services-HOWTO, updated: August 1998,

• WWW-HOWTO, updated: August 1999,

• WWW-mSQL-HOWTO, updated: September 1997,

• Wacom-Tablet-HOWTO, updated: Nov 1999,

• Wearable-HOWTO, updated: December 1999,

• XFree86-HOWTO, updated: May 1999,

• XFree86-Video-Timings-HOWTO, updated: January 2000,

• XWindow-User-HOWTO, updated: September 1999,

• 3-Button-Mouse, updated: May 1998,

• ADSL, updated: April 1999,

• ADSM-Backup, updated: January 1997,

• Advocacy, updated: May 1998, • Alsa-sound, updated: November 1999,

• Apache+SSL+PHP+fp, updated: July 1998,

• Automount, updated: April 1999,

• BTTV-Mini-HOWTO-0.3, updated: February 2000,

• Backup-With-MSDOS, updated: August 1997,

• Battery-Powered, updated: December 1997,

• BogoMips, updated: November 1999,

• Bridge+Firewall, updated: December 1997,

• Bridge, updated: September 1998,

• Bzip2, updated: August 1999,

• Cipe+Masq, updated: April 1999,

• Clock, updated: July 1999,

• Coffee, updated: January 1998,

• Commercial-Port-Advocacy, updated: December 1999,

• DPT-Hardware-RAID, updated: April 1999,

• Domain-Setup-mini-HOWTO, updated: and 2049,

• Ext2fs-Undeletion, updated: February 1999,

• FTP, updated: January 2000,

• Fax-Server, updated: November 1997,

• Firewall-Piercing, updated: November 1998,

• Firewall-and-ADSL-mini-HOWTO, updated:

• Font-Deuglification-mini-HOWTO, updated: November 1999,

• Home-Network-mini-HOWTO, updated: January 2000,

• IO-Port-Programming, updated: December 1997,

• IP-Subnetworking, updated: March 1997,

• ISP-Connectivity, updated: November 1997,

• Install-From-ZIP, updated: April 1998,

• LBX, updated: December 1997,

• LILO, updated: August 1998,

• Linux+DOS+Win95+OS2, updated: November 1997, • Linux+FreeBSD, updated: January 1999,

• Linux+NT-Loader, updated: September 1997,

• Loadlin+Win95, updated: March 1999,

• Loopback-Root-FS, updated: September 1999,

• Mac-Terminal, updated: November 1997,

• Mail-Queue, updated: September 1997,

• Mail2News, updated: July 1999,

• Multiboot-with-LILO, updated: March 1998,

• NCD-X-Terminal, updated: November 1999,

• NFS-Root, updated: August 1997,

• Netrom-Node, updated: October 1998,

• Netstation, updated: February 1998,

• News-Leafsite, updated: January 1998,

• Offline-Mailing, updated: June 1998,

• PLIP, updated: March 1998,

• Partition-Rescue-mini-HOWTO, updated: Jan 2000,

• Partition, updated: November 1997,

• Path, updated: November 1997,

• Programming-Languages, updated: January 2000,

• Public-Web-Browser, updated: January 1998,

• Qmail+MH, updated: March 1998,

• Quota, updated: August 1997,

• RCS, updated: August 1997,

• RPM+Slackware, updated: April 1998,

• RedHat-CD, updated: September 1999,

• Remote-Boot, updated: February 1999,

• Remote-X-Apps, updated: November 1999,

• Saving-Space, updated: April 1999,

• Secure-POP+SSH, updated: September 1998,

• Sendmail+UUCP, updated: May 1998, • Sendmail-Address-Rewrite, updated: May 1998,

• Software-RAID, updated: November 1998,

• Soundblaster-AWE, updated: January 1998,

• StarOffice, updated: November 1999,

• Swap-Space, updated: December 1999,

• TT-Debian, updated: January 2000,

• Term-Firewall, updated: July 1996,

• TkRat, updated: September 1999,

• Token-Ring, updated: January 1998,

• Ultra-DMA, updated: November 1999,

• Update, updated: September 1999,

• Upgrade, updated: June 1996,

• VAIO+Linux, updated: September 1998,

• VPN, updated: August 1997,

• Visual-Bell, updated: November 1997,

• WordPerfect, updated: August 1997,

• X-Big-Cursor, updated: August 1997,

• XFree86-XInside, updated: September 1997,

• Xterm-Title, updated: October 1999,

• ZIP-Drive, updated: January 1999,

• ZIP-Install, updated: January 1998,

3.2 Special HOWTOs

The High Availability HOWTO, by Harald Milz <hm@seneca.muc.de> is available at

http://metalab.unc.edu/pub/Linux/ALPHA/linux-ha/High-Availability-HOWTO.html.  It is not

included with the HOWTO collection because it relies on figures and cannot be distributed in all supported formats. The Graphics mini-HOWTO, by Michael J. Hammel <mjhammel@graphics-muse.org> is available at http://www.graphics-muse.org/linux/lgh.html. It is not included with the HOWTO collection because it needs to use a lot of images, which don’t translate to other output formats.

3.3 Unmaintained HOWTOs

There are a number of unmaintained documents.

These are kept around since old documentation is sometimes better than none. However, you should be aware that you are reading old documentation.

If the date tag is too old, please consider the HOWTO as unmaintained.

However, there are exceptions, such as old programs (for ex UUCP HOWTO) or not-so-hot topics (for ex AX-25 HOWTO).

4 Writing and Submitting a HOWTO

If you are interested in writing or maintaining a HOWTO, please get in touch with us FIRST!!! at

ldp-discuss@lists.linuxdoc.org.

If you want to maintain a unmaintained HOWTO, please also get in touch with the former author.

Here are a few guidelines that you should follow when writing a HOWTO:

• Try to use meaningful structure and organization, and write clearly. Remember that many of the people reading HOWTOs do not speak English as their first language.

• If you are writing a HOWTO, you must use the SGML-Tools package, available from http://www.sgmltools.org/, to format the HOWTO. This package allows us to produce HTML, LaTeX (for DVI and PostScript), plain text and DocBook sgml from a single source document, and was designed specifically for the HOWTOs. This also gives all of the HOWTOs a uniform look. It is very important that you format and review the output of the formatting in PostScript, plain text and HTML.

• Make sure that all of the information is correct. I can’t stress this enough. When in doubt, speculate, but make it clear that you’re only guessing.

• Make sure that you are covering the most recent version of the available software. Also, be sure to include full instructions on where software can be downloaded from (FTP site name, full pathname), and the current version number and release date of the software.

• Include an FAQ section at the end, if appropriate. Many HOWTO documents need an “FAQ” or “Common Problems” section to cover information which can’t be covered in the regular text.

• Use other HOWTOs a model! The SGML source to the HOWTOs is available on Linux FT-P sites. In addition, have a look at the LDP Style Guide <http://metalab.unc.edu/LDP/HOWTO/ LDP-Style-Guide.html>, the HOWTO-HOWTO, the LDP manifesto and the LDP license guide.

• Make sure that your name, email address, date, and a version number is near the beginning of the document. You could also include WWW addresses and a snail mail address if you want. The standard header is:

        Title
        Author's name and email address
        Version number and date

for example:

             The Linux HOWTO Index
             by Tim Bynum
             v2.10.29, 31 July 1997

• Lastly, be prepared to receive questions and comments about your writing. There are several hundreds of accesses to the HOWTO collection every day from around the world!

After you have written the HOWTO, mail the SGML source to us at ldp-submit@lists.linuxdoc.org.

It is important that you go through us when submitting a HOWTO, as we maintain the archives and need to keep track of what HOWTOs are being written and who is doing what.

Then, all you have to do is send me periodic updates whenever appropriate.

5 Copyright

Copyright (c) 1995 - 2000 by Tim Bynum. Copyright (c) 2000 by Guylhem Aznar.

Unless otherwise stated, Linux HOWTO documents are copyrighted by their respective authors. Linux HOWTO documents may be reproduced and distributed in whole or in part, in any medium physical or electronic, as long as this copyright notice is retained on all copies. Commercial redistribution is allowed and encouraged; however, the author would like to be notified of any such distributions.

All translations, derivative works, or aggregate works incorporating any Linux HOWTO documents must be covered under this copyright notice. That is, you may not produce a derivative work from a HOWTO and impose additional restrictions on its distribution. Exceptions to these rules may be granted under certain conditions; please contact the Linux HOWTO coordinator at the address given below.

In short, we wish to promote dissemination of this information through as many channels as possible. However, we do wish to retain copyright on the HOWTO documents, and would like to be notified of any plans to redistribute the HOWTOs.

If you have questions, please contact the Linux HOWTO coordinator, at linux-howto@metalab.unc.edu via email.